OpenStudy (anonymous):

factoring x^2+3x=8x-6

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

Fitst collect like terms. Can u do that?

OpenStudy (anonymous):

like what do you mean?

OpenStudy (anonymous):

\[x^{2}+3x=8x-6\]\[=x^{2}-5x+6=0\]\[=x^{2}-2x-3x+6=0\]\[=x(x-2)-3(x-2)=0\]\[=(x-3)(x-2)=0\]